OpenStudy (anonymous):

Assest the validity of the following statement: For all integrals m & n with m is opposite of -n & +n, and integral (cosm(t)cosn(t)dt)=0 from -pi to pi ???

myininaya (myininaya):

hey one sec i did this before

myininaya (myininaya):

let me see if i can find it

OpenStudy (anonymous):

yeaaaaaaay

OpenStudy (anonymous):

can't wait because i don't even understand the question

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wooww and how do you know that "we also know that thing?"

myininaya (myininaya):

i was just using identities we have \[\cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y)\] and \[\cos(x-y)=\cos(x)\cos(y)+\sin(x)\sin(y)\] adding these equations together we get \[\cos(x+y)+\cos(x-y)=2\cos(x)\cos(y)+0\] \[\cos(x+y)+\cos(x-y)=2\cos(x)\cos(y)\]

myininaya (myininaya):

multiply 1/2 on both sides and get \[\frac{1}{2}[\cos(x+y)+\cos(x-y)]=\cos(x)\cos(y)\]

myininaya (myininaya):

\[\cos(x)\cos(y)=\frac{1}{2}[\cos(x+y)+\cos(x-y)]\] \[\cos(m \theta) \cos(n \theta)=\frac{1}{2}[\cos(m \theta+n \theta)+\cos(m \theta- n \theta)]\] \[\cos(m \theta) \cos(n \theta)=\frac{1}{2}[\cos([m+n] \theta)+\cos([m-n] \theta)]\]

OpenStudy (anonymous):

so is this the answer ? or that one

myininaya (myininaya):

what do you mean?

myininaya (myininaya):

i went through the whole thing on the other thread i just put some more steps on this one because you asked so what do you mean which is the answer?

OpenStudy (anonymous):

ohh okk got it. thank youu!

myininaya (myininaya):

the whole thing from both threads is the answer this is a proof